OpenStudy (anonymous):

I GIVE GOLD MEDALS!!!!!!!! Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. A.) x-intercept: (0, −3), y-intercept: (0, 2) B.) x-intercept: (−3, 0), y-intercept: (2, 0) C.) x-intercept: (0, 2), y-intercept: (0, −3) D.) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (solomonzelman):

\[3x − 2y = 6.--->-2y=6-3x.-->y=-2+1/5x.-->y=1/5x-2\] Can you do it now?

OpenStudy (anonymous):

3x -2y = 6 so -2y = 6-3x so y = (6/-2) - (3/-2)x so y = -3 +(3/2)x this is in y = mx +b form so b = -3 which equals your y intercept. For the x intercept you can solve for x by setting y=0 and getting 0 = -3 + (3/2)x so 3=3/2x so x=2. when you write this out by points it should be written as (x,y) so your answer is x intercept = (2,0) and y intercept = (0, -3)